Determine the magnitude and direction of the electromagnetic force between the two charges 66 and -22 separated by the distance 45.



1
Expert's answer
2021-01-14T10:37:11-0500
"F=k\frac{q_1q_2}{r^2}\\
F=9\cdot10^9\frac{(66)(-22)}{(45)^2}=-6.45\cdot10^9N"

Negative sign means that the force is attractive.